OpenStudy (anonymous):

The area of a trapezoid is found using the formula 2003-01-03-00-00_files/i0080000.jpg, where 2003-01-03-00-00_files/i0080001.jpg and 2003-01-03-00-00_files/i0080002.jpg are the parallel sides of the trapezoid and h is the height of the trapezoid. What is the area, in square kilometers, of a trapezoid with bases 5 km and 7 km and height 3 km? A. 14.5 B. 18 C. 36 D. 52.5

Directrix (directrix):

The area of a trapezoid is 1/2*h*(b1 + b2) where h is the height, b1 and b2 are the two base lengths. A = 1/2*h*(b1 + b2) A = 1/2* 3 * (5 + 7) A = 1/2 * 3 * 12 A = ? @jkgjmgk